# Towards Real-Time Generation of Delay-Compensated Video Feeds for Outdoor Mobile Robot Teleoperation

### v1

@@ Abstract
Teleoperation is an important technology to enable supervisors to control agricultural robots remotely. However, environmental factors in dense crop rows and limitations in network infrastructure hinder the reliability of data streamed to teleoperators. These issues result in delayed and variable frame rate video feeds that often deviate significantly from the robot’s actual viewpoint. We propose a modular learning-based vision pipeline to generate delay-compensated images in real-time for supervisors. Our extensive offline evaluations demonstrate that our method generates more accurate images compared to state-of-the-art approaches in our setting. Additionally, we are one of the few works to evaluate a delay-compensation method in outdoor field environments with complex terrain on data from a real robot in real-time. Additional videos are provided at https://sites.google.com/illinois.edu/comp-teleop.
@@ I Introduction
While teleoperation is a necessary technology for these robots, we find through real-world testing that our organization’s existing teleoperation platform [^4] Details of the hardware behind the networking setup of the teleoperation system are left in a separate paper under review, and is out of scope. has several limitations, including severe delay in communication between the robot and supervisor, and intermittent transmission failures. These transmission issues are caused by inherent delay in sending information over a network with low bandwidth, and moisture in the crops causing signal fades in parts of the farm. Furthermore, the challenging under-canopy crop environment and unpredictable weather patterns lead to uneven terrain, causing large random deviations in consecutive camera poses. In extreme cases, severe delays may lead to undesired robot crashes, causing catastrophic task failure.
@@
Our primary contributions are summarized as follows: we (1) design an efficient modular pipeline for frame delay compensation [^1] Code will be released upon acceptance., (2) extensively compare our pipeline with ablations, classical image processing methods, and state-of-the-art learning-based image generation approaches on an offline crop dataset in diverse growth stages, and (3) showcase the real-time operation of our approach on real-world data by integrating our method into a ROS node.
@@ II-A Time Delay Compensation for Robot Teleoperation
Teleoperation is a primary mode of control in robotics for challenging tasks and environments where full automation is still actively under development, like space exploration [18], underwater operation [19], nuclear material handling [20], and more [21, 22, 23, 24]. In many of these applications, the human operator is expected to be far from the robot, introducing issues of limited bandwidth and perception latency [7]. In particular, high latency results in serious consequences where real-time responsiveness is critical [25]. Different methodologies have been investigated to mitigate the impact of delays on teleoperation, including devising a move-and-wait user strategy for tasks which allow quasi-static operations [26], and abstracting away low-level short-term control signals with high-level long-term user commands [27]. Another avenue of research for delay compensation is future frame prediction [11]. By incorporating the operator control commands, a simple method that uses sliding and zooming video transformation for prediction can achieve impressive performance gain in driving scenarios [13]. More recent efforts are harnessing neural networks to account for missing details from boundary disocclusions [15, 14, 17]. Note there are very few works on delay compensation for field robots due to the complexity of required networking infrastructure and the unstructured nature of outdoor environments [10].
@@ III-B Choosing a Scene Representation
One common approach to generating future frames conditioned on $I_{t-m:t}$, is to predict future $2$D pixel flows $F_{t+1:t+n}$, and iteratively applying the flows to $I_{t}$ [49, 50, 51]. Although effective in environments with smooth camera motions, flow-based methods fail in deployments where consecutive input frames appear far apart and have few correspondences, as is the case in choppy low-bandwidth networks like our setting.
@@ III-D Future Pose Prediction
where $\mu$ and $\eta$ are friction coefficients. Gasparino et al[62] predict these coefficients with a learned neural network. To improve runtime efficiency and reduce memory usage, we set $\mu=\eta=1$ during evaluation, effectively simplifying the kinematics to the Dubins’ car model [63].
@@ III-E Rendering with Pulsar
where $K_{w}$ and $K_{f}$ are the camera’s sensor width and focal length respectively. Intuitively, points closer to the camera are given smaller radii during rendering. Unlike the sampling procedure of early neural radiance field approaches [64], Pulsar only stores memory for occupied regions in space, enabling efficient computation of the rendered color for a pixel. In particular, Pulsar uses a softmax blending function to weight overlapping points along a ray when rendering the projected image (Eq. 1 from Lassner et al[39]). Increasing the $\gamma$ parameter in this function enables us to raise the transparency of nearby points. Finally, and arguably most important for real-world deployment, Pulsar is integrated into PyTorch [65], eliminating the bottleneck of transferring DAv2 predictions or point clouds between libraries or devices — required for other renderers [66, 67], and it is implemented with specialized CUDA kernels that allow for real-time rendering of images on our hardware.
@@ IV-B Results
As an ablation study, we train an additional inpainting model with the same architecture using MS-SSIM as its loss function, following recent findings from Shi et al[70], who report the superiority of structural similarity index measure-based losses for image generation. Qualitatively, we find the raw predictions output by ResNet-MS-SSIM are sharper than the model trained with L1 loss in the regions without disocclusions in the reprojected image. However, once we use the raw prediction to fill in holes from the reprojection, we see the painted regions contrast heavily with the remainder of the image. On the other hand, while L1 loss leads to blurrier raw predictions, the predictions for disoccluded patches in reprojected image blend in smoother, resulting in a higher PSNR. Thus, we use the ResNet model trained with L1 loss as our inpainting model. Examples of model predictions are provided in Figure 4.
@@
We present results on the accuracy of model predictions across each test video and different time delays in Table III and example generations are provided in the supplentary video. Generally, we find all methods perform best on early stage sequences compared to middle and late stage images due to fewer occlusions. Similarly, as expected, larger delays lead to worse generations. However, on average across the board, our method outperforms the others in all three metrics. Particularly, C$+$S generates cropped and resized images which have high perceptual similarity to the original scene, but in fact align poorly with the true image. Iterative approaches like SRVP and DMVFN incrementally produce more blurred, incomprehensible outputs as intermediate errors compound. Finally, we find SynSin is unable to learn accurate enough depth or CNN features to decode future states accurately. SRVP realizes the closest results to ours in late stage PSNR, but it requires training over five days on two NVIDIA A100 GPUs with $256\times 256$ resolution images, whereas finetuning DAv2 and training our inpainting model each took one day on half the compute with the full resolution images. DMVFN similarly achieves comparable MS-SSIM, but its blurry generations result in poor LPIPS. However, it is worth noting that C$+$S, SynSin, and our model predicts future frames assuming the world is static, resulting in inaccurate results when wind or the robot itself moves crops. We also find all methods perform poorly when conditioning generations on an occluded frame, leading us to develop real-time occlusion filters in future work.
@@ V Real-Time Experiments
We also develop a ROS node to deploy our delay compensation method in real-time on TerraSentia rosbags. A block diagram of the node is shown in Figure 2. To test the quality of the real-time compensated video feed, we emulate the conditions of varying network settings by skipping $5$ and $10$ frames (effectively requiring to compensate for $6$ and $3$ FPS videos respectively, from a $30$ Hz stream), and applying $250$ and $500$ ms delays to real-world rosbags in different growth stages. Resulting videos are provided in the supplementary material. We qualitatively find our model is capable of compensating for different frame rates and delays. However, noisy odometry measurements and poor predictions from our simplistic kinematic model lead to undesired jumps in the generated video under worse conditions.

### v2

@@ Abstract
Teleoperation is an important technology to enable supervisors to control agricultural robots remotely. However, environmental factors in dense crop rows and limitations in network infrastructure hinder the reliability of data streamed to teleoperators. These issues result in delayed and variable frame rate video feeds that often deviate significantly from the robot’s actual viewpoint. We propose a modular learning-based vision pipeline to generate delay-compensated images in real-time for supervisors. Our extensive offline evaluations demonstrate that our method generates more accurate images compared to state-of-the-art approaches in our setting. Additionally, ours is one of the few works to evaluate a delay-compensation method in outdoor field environments with complex terrain on data from a real robot in real-time. Resulting videos and code are provided at https://sites.google.com/illinois.edu/comp-teleop.
@@ I Introduction
While teleoperation is a necessary technology for these robots, we find through real-world testing that our organization’s existing teleoperation platform [^4] Details of the hardware behind the networking setup of the teleoperation system are discussed by Sie et al[7], and is out of the scope of this work. has several limitations, including severe delay in communication between the robot and supervisor, and intermittent transmission failures [7]. These transmission issues are caused by inherent delay in sending information over a network with low bandwidth, and moisture in the crops causing signal fades in parts of the farm. Furthermore, the challenging under-canopy crop environment and unpredictable weather patterns lead to uneven terrain, causing large random deviations in consecutive camera poses. In extreme cases, severe delays may lead to undesired robot crashes, causing catastrophic task failure.
@@
Our primary contributions are summarized as follows: we (1) design an efficient modular pipeline for frame delay compensation, (2) extensively compare our pipeline with ablations, classical image processing methods, and state-of-the-art learning-based image generation approaches on an offline crop dataset in diverse growth stages, and (3) showcase the real-time operation of our approach on real-world data by integrating our method into a ROS node.
@@ II-A Time Delay Compensation for Robot Teleoperation
Teleoperation is a primary mode of control in robotics for challenging tasks and environments where full automation is still actively under development, like space exploration [19], underwater operation [20], nuclear material handling [21], and more [22, 23, 24, 25]. In many of these applications, the human operator is expected to be far from the robot, introducing issues of limited bandwidth and perception latency [8]. In particular, high latency results in serious consequences where real-time responsiveness is critical [26]. Different methodologies have been investigated to mitigate the impact of delays on teleoperation, including devising a move-and-wait user strategy for tasks which allow quasi-static operations [27], and abstracting away low-level short-term control signals with high-level long-term user commands [28]. Another avenue of research for delay compensation is future frame prediction [12]. By incorporating the operator control commands, a simple method that uses sliding and zooming video transformation for prediction can achieve impressive performance gain in driving scenarios [14]. More recent efforts are harnessing neural networks to account for missing details from boundary disocclusions [16, 15, 18]. Note that there are very few works on delay compensation for field robots due to the complexity of required networking infrastructure and the unstructured nature of outdoor environments [11].
@@ III-B Choosing a Scene Representation
One common approach to generating future frames conditioned on $I_{t-m:t}$, is to predict future $2$D pixel flows $F_{t+1:t+n}$, and iteratively apply each flow to $I_{t}$ [50, 51, 52]. Although effective in environments with smooth camera motions, flow-based methods fail in deployments where consecutive input frames appear far apart and have few correspondences, as is the case in choppy low-bandwidth networks like our setting.
@@ III-D Future Pose Prediction
where $\mu$ and $\eta$ are friction coefficients. Gasparino et al[63] predict these coefficients with a learned neural network. To improve runtime efficiency and reduce memory usage, we set $\mu=\eta=1$ during evaluation, effectively simplifying the kinematics to the Dubins’ car model [64]. We note, however, that $\mu$ and $\eta$ may be tuned or adaptively chosen in real-time for the specific terrain.
@@ III-E Rendering with Pulsar
where $K_{w}$ and $K_{f}$ are the camera’s sensor width and focal length respectively. Intuitively, points closer to the camera are given smaller radii during rendering. Unlike the sampling procedure of early neural radiance field approaches [65], Pulsar only stores memory for occupied regions in space, enabling efficient computation of the rendered color for a pixel. In particular, Pulsar uses a softmax blending function to weight overlapping points along a ray when rendering the projected image (Eq. 1 from Lassner et al[40]). Increasing the $\gamma$ parameter in this function enables us to raise the transparency of nearby points. Finally, and arguably most important for real-world deployment, Pulsar is integrated into PyTorch [66], eliminating the bottleneck of transferring DAv2 predictions or point clouds between libraries or devices — required for other renderers [67, 68] — and it is implemented with specialized CUDA kernels that allow for real-time rendering of images on our hardware.
@@ IV-B Results
As an ablation study, we train an additional inpainting model with the same architecture using MS-SSIM as its loss function, following recent findings from Shi et al[71], who report the superiority of structural similarity index measure-based losses for image generation. Qualitatively, we find the raw predictions output by ResNet-MS-SSIM are sharper than the model trained with L1 loss in the regions without disocclusions in the reprojected image. However, once we use the raw prediction to fill in holes from the reprojection, we see the painted regions contrast heavily with the remainder of the image. On the other hand, while L1 loss leads to blurrier raw predictions, the predictions for disoccluded patches in the reprojected image blend in smoother, resulting in a higher PSNR. Thus, we use the ResNet model trained with L1 loss as our inpainting model. Examples of model predictions are provided in Figure 4.
@@
We present results on the accuracy of model predictions across each test video and different time delays in Table III and example generations are provided in the supplentary video. Generally, we find all methods perform best on early stage sequences compared to middle and late stage images due to fewer occlusions. Similarly, as expected, larger delays lead to worse generations. However, on average across the board, our method outperforms the others in all three metrics. Particularly, C$+$S generates cropped and resized images which have high perceptual similarity to the original scene, but in fact align poorly with the true image. Iterative approaches like SRVP and DMVFN incrementally produce more blurred, incomprehensible outputs as intermediate errors compound. Finally, we find SynSin is unable to learn accurate enough depth or CNN features to decode future states accurately. SRVP realizes the closest results to ours in late stage PSNR, but it requires training over five days on two NVIDIA A100 GPUs with $256\times 256$ resolution images, whereas finetuning DAv2 and training our inpainting model each took one day on half the compute with the full resolution images. DMVFN similarly achieves comparable MS-SSIM, but its blurry generations result in poor LPIPS. However, it is worth noting that C$+$S, SynSin, and our model predict future frames assuming the world is static, resulting in inaccurate results when wind or the robot itself moves crops. We also find all methods perform poorly when conditioning generations on an occluded frame, leading us to develop real-time occlusion filters in future work.
@@ V Real-Time Experiments
We also develop a ROS node to deploy our delay compensation method in real-time on TerraSentia rosbags. A block diagram of the node is shown in Figure 2. To test the quality of the real-time compensated video feed, we emulate the conditions of varying network settings by skipping $5$ and $10$ frames (effectively requiring to compensate for $6$ and $3$ FPS videos respectively, from a $30$ Hz stream), and applying $250$ and $500$ ms delays to real-world rosbags in different growth stages on straight-path and turning trajectories. Results are in the supplementary material and website. We qualitatively find our model is capable of compensating for different frame rates and delays. However, noisy odometry and poor predictions from our simplistic kinematic model lead to undesired jumps in the generated video under worse conditions.